OpenStudy (anonymous):

Let f(x) = x^4 – 8x^3 + 16x^2 – 19 and g(x) = x – 5. Find f(x)/g(x).

OpenStudy (anonymous):

\[\frac{f(x)}{g(x)}=\frac{x^4-8x^2+16x^2 -19}{x-5}\]

jimthompson5910 (jim_thompson5910):

use polynomial long division luckily wolfram alpha does that quite well http://www.wolframalpha.com/input/?i= \frac{x^4-8x^2%2B16x^2+-19}{x-5}

jimthompson5910 (jim_thompson5910):

you'll have to copy and paste that link since clicking the link won't work

OpenStudy (anonymous):

mkay thanks!

jimthompson5910 (jim_thompson5910):

you're welcome